OpenStudy (rissa):

A line passes through (–10, –6) and (–6, 1). a. Write an equation for the line in point-slope form. b. Rewrite the equation in standard form using integers.

OpenStudy (anonymous):

\[y-y_{1}=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}(x-x_{1})\] \[y+6=\frac{-6-1}{-10+6}(x+10)\] \[y+6=\frac{7}{4}(x+10)\] that is point slope form